placed in a uniform magnetic field; the wire makes an angle of Owith the (9) A straight wire is magnetic field. The wire carries a current i, and the magnetic field strength is T. Calculate the force on a segment of length of this wire. Show the direction of the force in a sketch, which also indicates the direction of the field and the magnetic field. [20 pts]
Answer #1

\vec{F}=i\vec{L}\times\vec{T}

where L vector is in the direction of current.

F=iLT\sin\theta

The direction of force is out of the page.